What is the answer to (64 divided by 8 squared plus 9 times 3 plus 7 squared) squared? Thank you so very much!
Answer by stanbon(75887)

What is the answer to (64 divided by 8 squared plus 9 times 3 plus 7 squared) squared?
------------------
I'll take a stab at it; but it is ambiguous.
(64/8^2 + 9*3 + 7^2)^2
-----------------------
= (1 + 27 + 49)^2
------
= (77)^2
----
= 77*77
---------
= 5929
============
Cheers,
Stan H.
============